# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@ 5. Empirical Evaluation
We now evaluate the effectiveness of NSDEs, as a DP-learner, in mitigating membership inference attacks. Our evaluation is conducted on the CIFAR-10 and CIFAR-100 datasets. Our NSDE models (SDENet) use architectural configurations similar to those of ODENet, consisting of three groups of ODEBlocks with filters with 16, 32, and 64 dimensions, respectively, widened by a factor of 2. In SDENet, stochasticity is introduced by employing an SDE solver within the ODEBlocks, utilizing a stochastic fourth-order Runge-Kutta (RK4) integration method. SDENet also incorporates a hyper-parameter ($\sigma$) to control noise intensity, influenced by factors such as time, step size, and the desired level of stochasticity.
5.1. Effectiveness of NSDEs
@@
Methodology. Here we conduct three comparisons: (1) We compare NSDEs with ResNet-14 to assess whether stochasticity reduces the empirical membership risks. (2) We compare NSDEs with empirical defenses discussed in Sec 2, which lack provable privacy guarantees. (3) We compare NSDEs with “private” ResNet-14 trained with DP-SGD. In (2), we consider four defenses studied in prior work (Kaya et al., 2020; Jia et al., 2019; Li et al., 2021; Ponomareva et al., 2023): $\ell^{1}$/$\ell^{2}$-regularization, MMD-Mixup, and MemGuard. For $\ell^{1}$/$\ell^{2}$-regularization, we set the penalty parameter $\lambda$ to $10^{-5}$. We use the most robust version of MemGuard as described in Choquette et al. (Choquette-Choo et al., 2021), which allows for arbitrary adjustments to the confidence vector while preserving the model’s predicted label. For MMD-Mixup, we adopt the implementation from the study by Li et al. (Li et al., 2021), and set the penalty parameter $\lambda$ to 5×$\times 10^{-4}$. When training with DP-SGD, we set the privacy budget $\epsilon$ to 8, a common choice for CIFAR-10/-100 in prior work (Ponomareva et al., 2023). Note that we use the popular Opacus library [^3] Opacus: https://github.com/pytorch/opacus/ to implement privacy accounting in NSDE training and to compute the total privacy leakage ($\varepsilon$).
@@
Figure 3. Comparing the effectiveness of different defenses, in CIFAR-10. We use LiRA to assess their membership risks. All defenses, except for SDENets, are applied to ResNet-14.
Figure 4. Impact of privacy hyper-parameters on NSDE training in CIFAR-10. We analyze the total privacy leakage ($\varepsilon$) in the left two plots and the privacy-utility tradeoff in the right two plots, while varying key hyper-parameters.
@@
Defense effectiveness. Figure 3 summarizes our results against LiRA, the strongest-known membership inference attack by Carlini et al (Carlini et al., 2022). Overall, we first demonstrate that NSDEs effectively mitigate membership risks. Compared to ResNet-14, a feedforward model that achieves 3.96–10.59% TPR at both 0.1% and 1% FPR, NSDEs with $\sigma\!=\!2$ demonstrate approximately two orders of magnitude lower TPRs (0.14–1.24%). We also find that, compared to existing defenses, NSDEs are significantly more effective in reducing membership risks. Compared to the four empirical defenses—$\ell^{1}$/$\ell^{2}$-regularization, MMD-Mixup and MemGuard—NSDEs with $\sigma\!=\!2$ achieves 2–8$\times$ lower TPR at both 0.1% and 1% FPR. Notably, NSDEs achieve effectiveness comparable to ResNet-14 trained with DP-SGD ($\varepsilon\!=\!8$). Moreover, it is interesting to observe that NSDEs with $\sigma\!=\!0$, the non-private variant of NSDEs, show lower TPRs compared to defended models.
@@
NSDEs with $\sigma\!=\!0$ achieve a test accuracy of 83.9%, which, although lower than ResNet-14 at 86.0%, significantly narrows the generalization gap from 12.2% to 5.2%. This reduction results in a markedly lower TPR at 0.1% FPR of 0.4%, compared to 4.0% for ResNet-14. NSDEs with $\sigma\!=\!2$ attain a test accuracy of 81.9% and virtually eliminate the generalization gap (-0.02%), further lowering the TPR at 0.1% FPR to 0.1%. When compared to $\ell^{1}$ and $\ell^{2}$-regularization, NSDEs with $\sigma\!=\!2$ not only provide a modest reduction in test accuracy by 4.6%–4.7%, but also enhance defenses, achieving a TPR at 0.1% FPR of 0.14% versus 1.1% for $\ell^{1}$ and 2.4% for $\ell^{2}$). While DP-SGD with $\varepsilon\!=\!8$ realizes a smallest generalization gap, it incurs significant sacrifices in test accuracy, ranging from 7.5% and 9.5% when compared to NSDEs ($\sigma\!=\!0$ and $\sigma\!=\!2$) Consequently, NSDEs demonstrates an improved overall performance in terms of both model training (higher test accuracy) and defense against membership inference attacks, making it a more balanced and practical approach. Please refer to Appendix C for more details.
@@
To conduct an in-depth analysis, we present the membership risks of NSDEs and NODEs against five existing attacks in Table 2. The attack success, measured across four metrics, is 4–10$\times$ lower in NSDEs compared to NODEs. Particularly against LiRA, NSDEs show a TPR of 0.14–1.24% at both 0.1% and 1% FPRs, whereas 1.0–4.22% TPR at the same FPRs in NODEs. To understand where the resilience of NSDEs against these attacks comes, we compare the generalization gap. In NODEs, the gap is 9.9–15.8% points for CIFAR-10 and CIFAR-100, while in NSDEs, it is 0.0–5.9% points. This result implies that the reduction in membership risks in NSDEs is partly due to a significant reduction in overfitting.
5.2. In-depth Analysis of Privacy of NSDEs
Figure 5. Comparing the empirical membership risks of NSDEs with the theoretical bound by DP-SGD. We use LiRA to measure the membership risks and the theoretical bound when $\varepsilon\!=\!1$ is computed by the method presented in (Carlini et al., 2022). NSDE with $\sigma\!=\!2$ is used to compare, and that with $\sigma\!=\!0$ is its non-private version. The shaded areas give the 68% (darker) and 95% (lighter) confidence intervals.
@@ 5.2. In-depth Analysis of Privacy of NSDEs
We next analyze the privacy guarantee ($\varepsilon$) provided by NSDEs. Because we theoretically show that NSDEs are DP-learners (with the privacy guarantee equivalent to those of DP-SGD), it is important to understand empirically whether their privacy interactions behave similarly to observations from prior work (Jayaraman and Evans, 2019). We first compare the privacy leakage of NSDEs with the worst-case leakage bound established by DP-SGD. We also examine the impact of privacy hyper-parameters—such as $\delta$ and the choice of privacy accounting algorithms—on the guarantee $\varepsilon$ and the model utility.
@@
Comparison to the theoretical DP-SGD bound. Figure 5 compares the upper-bound of DP-SGD (when $\varepsilon\!=\!1$) with the LiRA success on NSDEs with $\sigma\!=\!2$. Note that training NSDEs with $\sigma\!=\!2$ while targeting $\varepsilon\!=\!1$ completely destroys the model utility, making any meaningful LiRA success unattainable. We therefore target $\varepsilon\!=\!8$ for NSDEs with $\sigma\!=\!2$. This does not compromise the generality of our results, because even against NSDEs with weaker privacy guarantees, the empirical attack (LiRA) does not exceed the DP-SGD bound. We first observe that the attack success of LiRA on NSDEs with $\sigma\!=\!2$ remains within the worst-case leakage bound established by DP-SGD. However, when the diffusion mechanism is absent ($\sigma\!=\!0$), the attack success often goes beyond the DP-SGD bound, particularly in lower FPR regions (e.g., at 0.1–1% FPR), indicating that privacy is not guaranteed.
We also investigate the impact of privacy hyper-parameters on the guarantee ($\varepsilon$) and the privacy-utility tradeoff.
@@
Privacy leakage $\varepsilon$ over training. The left two figures in Figure 4 illustrate the total privacy leakage $\varepsilon$ over training. Both figures show that privacy leakage increases with training epochs, as expected, but the growth rate decreases over time. This aligns with empirical findings from prior work on training models with DP-SGD (Abadi et al., 2016).
@@
The leftmost figure presents the privacy leakage computed using three different privacy accountants: Gaussian Differential Privacy (GDP), Renyi Differential Privacy (RDP), and Privacy loss Random Variables (PRVs). GDP accountants consistently yield higher estimates of privacy leakage compared to the other two methods. This is partly because GDP relies on Gaussian approximations, which often deviate from the true privacy loss distribution. In contrast, precise accounting methods like RDP or PRVs, which enable tracking the exact or near-exact cumulative privacy loss, provide tighter privacy bounds. Due to their cumulative nature, we also observe that the differences among the three accountants progressively widens as the number of training epochs increases.
The figure second from the left shows the privacy leakage computed by the GDP accountant across various $\delta$ values, ranging from $1\times 10^{-2}$ to $1\times 10^{-6}$. Our observations align with the literature: as $\delta$ increases, the total privacy leakage ($\varepsilon$) also increases. This is because $\delta$ represents the probability of catastrophic events where the mechanism may fail to provide privacy guarantees, such as a database breach caused by physical attacks. A larger $\delta$ necessitates a relaxation of the privacy requirement, effectively allowing for a slightly higher probability of privacy violations.
@@
Figure 6. Membership risks of NSDEs trained with different configurations. (Top row, left to right) We illustrate the impact of the number of blocks ($n$), the level of stochasticity ($k$), and the step-size ($s$). (Bottom row, left to right) We show the impact of the integration time $T$ and the noise intensity $\sigma$. We use LiRA to assess the membership risks. The results are in CIFAR-10.
@@
Privacy-utility trade-off. The right two figures in Figure 4 illustrate the test accuracy of NSDEs trained with different privacy budgets ($\varepsilon$). We record the test accuracy and privacy leakage at each training epoch for the 16 models used in our evaluation, and we plot them. The second figure from the right shows the trade-offs across different failure probabilities ($\delta$). As $\delta$ increases, the test accuracy of NSDEs increases under the same privacy budget. This is because a larger $\delta$ permits a higher likelihood of privacy leakage. Surprisingly, NSDEs achieve a test accuracy of 80% with $\varepsilon$ = 2 and $\delta$ = $10^{-5}$ under GDP, which demonstrates their ability to maintain high utility even at relatively low privacy budgets. The rightmost figure demonstrates that all three accountants produce similar estimates of privacy leakage and test accuracy. However it is important to note the variability in test accuracy: even under the same privacy budget, e.g., $\varepsilon\!=\!2$, test accuracy ranges from 65–82%. We attribute this variance to the factors like randomness introduced by the diffusion term or the training data stocasticity. This suggests that in practice, to train NSDEs with high utility we may need to train the same model multiple times, as shown in prior work (Ponomareva et al., 2023).
5.3. Impact of NSDE Configurations
@@
The top-left panel in Figure 6 illustrates our results. Membership risks associated with NSDE models are consistent when $n\!=\!1$ or $n\!=\!2$. But employing three SDE blocks reduces the risks by a factor of 2, as measured by TPRs at 0.1–1% FPR. We attribute this reduction to poor generalization when $n\!=\!3$. For instance, in CIFAR-10, while NSDEs with $n\in\{1,2\}$ achieve test accuracy of 81.9–82.4%, models with $n\!=\!3$ show significantly lower test accuracy of 74.6%. Interestingly, unlike NODEs, augmenting the number of SDE blocks in NSDEs leads to a corresponding reduction in membership risks.
@@
The top-middle panel in Figure 6 illustrates our findings. We show that NSDEs with varying $k$ values effectively mitigate membership risks, as measured by LiRA. With $k$ values up to 0.5, NSDEs exhibit the same level of effectiveness. But when $k$ reaches to 0.6, it completely mitigates the membership risks, bringing the ROC curve below that of a random classifier (a diagonal line). We also examine how each choice of $k$ affects model performance. Increasing the level of stochasticity results in a decrease in model utility. On CIFAR-10, the test accuracy drops from 82.6% to 57.1% as $k$ increases from 0.3 to 0.6. This result implies that the state-of-the-art attack (LiRA) is yet weak so that one can completely mitigate with $k$ that does not significantly degrade model utility.
@@
The top-right panel in Figure 6 illustrates our findings. Overall, we find no significant interaction between the integration time interval ($T$) and membership risks. All NSDE models consistently show low membership risks. But theoretically, decreasing $T$ will increase the noise intensity ($\sigma$), leading to a smaller privacy leakage ($\varepsilon$) with reduced model utility and membership risks.
@@
The bottom-left panel in Figure 6 summarizes our findings. Overall, we observe only marginal variations in membership risks caused by changes in step size ($1/s$). On CIFAR-10, the membership risks observed at $s\!=\!8$ or $16$ are comparable. However, we find a reduction of the risks with $s\!=\!32$, likely because the model lacks learning detailed dynamics from the training data. Our analysis indicates that this reduction comes at the cost of accuracy. A step size of $32$ limits NSDEs effectiveness in modeling underlying dynamics, and in consequence, they experience substantial performance degradation. NSDEs at this step size achieve an average accuracy of 35.9%, which is 46.5–46.1% lower than that achieved with $s$ of 8 or 16.
@@
The bottom-right panel in Figure 6 illustrates our results. Overall, we observe that all the noise intensity values we examine effectively mitigate membership risks, consistent with prior work (Carlini et al., 2022), which shows that even the strongest membership inference attacks can be easily countered by a weak privacy guarantee provided by DP-SGD. We also demonstrate that as the noise intensity gets stronger, NSDE’s effectiveness increases. However, it is important how we configure the noise intensity. For instance, setting a lower noise intensity, such as 1.6, by employing a larger step-size (e.g., 32), may appear to reduce membership riusks in the figure. But this reduction is primarily due to a substantial drop in model utility, as shown in our analysis of different step-sizes.
5.4. NSDE as a Drop-in Replacement Module

### v2

@@ 5 Empirical Evaluation
We now empirically evaluate the effectiveness of NSDEs, as a DP-learner, in mitigating membership inference attacks. We run our evaluation on the CIFAR-10 and CIFAR-100 datasets, where membership inference attacks are shown to be most effective. Our NSDE models (SDENet) adopt architectural configurations similar to ODENet, comprising three groups of ODEBlocks with 16, 32, and 64 dimensional filters, respectively, widened by a factor of 2. In SDENet, stochasticity is introduced by employing an SDE solver within the ODEBlocks, utilizing a stochastic fourth-order Runge-Kutta (RK4) integration method. SDENet also incorporates a hyper-parameter ($\sigma$) to control noise intensity, influenced by factors such as time, step size, and the desired level of stochasticity.
5.1 Effectiveness of NSDEs
@@
We conduct three comparisons: (1) NSDEs with ResNet-14 to assess whether stochasticity reduces the empirical membership risks. (2) NSDEs with heuristic defenses discussed in §2, which lack provable privacy guarantees. (3) NSDEs with “private” ResNet-14 trained with DP-SGD. In (2), we consider four defenses studied in prior work [24, 36, 37, 47]: $\ell^{1}$/$\ell^{2}$-regularization, MMD-Mixup, and MemGuard. For $\ell^{1}$/$\ell^{2}$-regularization, we set the penalty parameter $\lambda$ to $10^{-5}$. We use the most robust version of MemGuard as described in Choquette et al. [48], which allows for arbitrary adjustments to the confidence vector while preserving the model’s predicted label. For MMD-Mixup, we adopt the implementation from the study by Li et al. [37], and set the penalty parameter $\lambda$ to 5×$\times 10^{-4}$. In (3), when training with DP-SGD, we set the privacy budget $\epsilon$ to 8, a common choice for CIFAR-10 and 100 in prior work [47]. We use the popular Opacus library [^3] Opacus: https://github.com/pytorch/opacus/ to implement privacy accounting in NSDE training and to compute the total privacy leakage ($\varepsilon$).
(1) Mitigating membership inference attacks. We first compare the membership inference risks of NSDEs and NODEs against six existing attacks. Table III shows our results. Across four evaluation metrics, NSDE reduce attack success by 4–10$\times$ compared to NODEs. Particularly against the LiRA attack, NSDEs achieve a TPR of 0.14–1.24% at both 0.1% and 1% FPRs, while NODEs exhibit a higher TPR of 1.0–4.22% under the same conditions (FPRs).
@@
Figure 3: Comparing the effectiveness of different defenses, in CIFAR-10. We use LiRA to assess their membership risks. All defenses, except for SDENets, are applied to ResNet-14.
@@
(2) Comparing with heuristic defenses. Figure 3 summarizes our results against LiRA, the membership inference attack where we observe the most significant reduction in attack success. Overall, we first demonstrate that NSDEs effectively mitigate membership inference risks. Compared to ResNet-14, a feedforward model that yields 3.96–10.57% TPR at both 0.1% and 1% FPR, NSDEs with $\sigma\!=\!2$ achieve approximately two orders of magnitude lower TPRs (0.14–1.24%). We also find that NSDEs significantly outperform heuristic defenses in reducing membership inference risks. Compared to the four empirical defenses—$\ell^{1}$/$\ell^{2}$-regularization, MMD-Mixup and MemGuard—NSDEs with $\sigma\!=\!2$ achieves 2–8$\times$ lower TPRs at 0.1–1% FPRs.
@@
(3) Comparison with private models constructed via DP-SGD. NSDEs match the effectiveness of ResNet-14 trained with DP-SGD at $\varepsilon\!=\!8$. It is interesting to observe that even the non-private variant of NSDEs (i.e., $\sigma\!=\!0$) achieve lower TPRs than these defended models. NSDEs with $\sigma\!=\!0$ achieve a test accuracy of 83.9%, which, though slightly lower than ResNet-14’s 86.0%, significantly narrows the generalization gap from 12.2% to 5.2%. This narrowing leads to a markedly lower TPR at 0.1% FPR: 0.4% for NSDEs versus 4.0% for ResNet-14. NSDEs with $\sigma\!=\!2$ attain a test accuracy of 81.9% and nearly eliminate the generalization gap (-0.02%), further reducing the TPR to 0.1% at 0.1% FPR. Compared to $\ell^{1}$ and $\ell^{2}$-regularization, NSDEs with $\sigma\!=\!2$ not only provide a modest reduction in test accuracy by 4.6%–4.7%, but also enhance defenses, achieving a TPR at 0.1% FPR of 0.14% versus 1.1% for $\ell^{1}$ and 2.4% for $\ell^{2}$). While DP-SGD with $\varepsilon\!=\!8$ realizes a smallest generalization gap, it significant compromises test accuracy, dropping 7.5–9.5% compared to NSDEs ($\sigma\!=\!0$ and $\sigma\!=\!2$). Consequently, NSDEs demonstrates an improved overall performance in terms of both model training (higher test accuracy) and defense against membership inference attacks, making it a more balanced and practical approach. Please refer to Appendix for further details.
Figure 4: Impact of privacy hyper-parameters on NSDE training in CIFAR-10. We analyze the total privacy leakage ($\varepsilon$) in the left two plots and the privacy-utility tradeoff in the right two plots, while varying key hyper-parameters.
@@
Moreover, to understand where the resilience of NSDEs against these attacks comes, we compare the generalization gap in Table IV. In NODEs, the gap is 9.9–29.89% for CIFAR-10 and CIFAR-100, while in NSDEs, it is 0.0–5.9%. This result implies that the reduction in membership risks in NSDEs is partly due to a significant reduction in overfitting.
TABLE IV: Generalization gap, measured for both NODE models and NSDE models across four benchmark datasets.
Task F-M C-10 C-100 T-I Model NODE NSDE NODE NSDE NODE NSDE NODE NSDE Train 95.5% 96.2% 94.3% 81.9% 82.1% 56.5% 55.9% 39.9% Test 90.0% 89.2% 84.4% 81.9% 52.2% 50.6% 40.5% 38.1% $\Delta$ 5.5% 7.0% 9.9% 0.0% 29.9% 5.9% 15.4% 1.8%
5.2 In-depth Analysis of Privacy of NSDEs
@@
We next analyze the privacy guarantee ($\varepsilon$) provided by NSDEs. Because we theoretically show that NSDEs are DP-learners (with the privacy guarantee equivalent to those of DP-SGD), it is important to understand empirically whether their privacy interactions behave similarly to observations from prior work [49]. We first compare the privacy leakage of NSDEs with the worst-case leakage bound established by DP-SGD. We also examine the impact of privacy hyper-parameters—such as $\delta$ and the choice of privacy accounting algorithms—on the guarantee $\varepsilon$ and the model utility.
Figure 5: Comparing the empirical membership risks of NSDEs with the theoretical bound by DP-SGD. We use LiRA to measure the membership risks and the theoretical bound when $\varepsilon\!=\!1$ is computed by the method presented in [6]. NSDE with $\sigma\!=\!2$ is used to compare, and that with $\sigma\!=\!0$ is its non-private version. The shaded areas give the 68% (darker) and 95% (lighter) confidence intervals.
@@
Comparison to the theoretical DP-SGD bound. Figure 5 compares the empirical success of LiRA against NSDEs with $\sigma\!=\!2$ to the theoretical upper-bound of DP-SGD (when $\varepsilon\!=\!1$). Note that training NSDEs with $\sigma\!=\!2$ while targeting $\varepsilon\!=\!1$ completely destroys the model utility, making any meaningful LiRA success unattainable. To maintain utility, we target $\varepsilon\!=\!8$ in our experiments with NSDEs. This does not compromise the generality of our results, because even against NSDEs with weaker privacy guarantees, the empirical attack (LiRA) does not exceed the DP-SGD bound. We first observe that the attack success of LiRA on NSDEs with $\sigma\!=\!2$ remains within the worst-case leakage bound established by DP-SGD. However, when the diffusion mechanism is removed ($\sigma\!=\!0$), LiRA’s attack success often goes beyond the DP-SGD bound, especially at lower FPR regions (e.g., at 0.1–1% FPR), highlighting the absence of formal privacy guarantees in non-stochastic variants.
@@
Privacy leakage $\varepsilon$ over training. The left two plots in Figure 4 illustrate the total privacy leakage ($\varepsilon$) during NSDE training. As expected, both plots show that privacy leakage increases with training epochs. However, the growth rate decreases over time—consistent with empirical findings from prior work on training models with DP-SGD [23].
@@
The leftmost plot compares privacy leakage estimates produced by three different accounting methods: Gaussian Differential Privacy (GDP), Renyi Differential Privacy (RDP), and Privacy loss Random Variables (PRVs). Among them, GDP consistently yield higher estimates of privacy leakage compared to the other two methods. This is partly because GDP relies on Gaussian approximations, which often deviate from the true privacy loss distribution. In contrast, precise accounting methods like RDP or PRVs, which enable tracking the exact or near-exact cumulative privacy loss, provide tighter privacy bounds. Moreover, because privacy loss accumulates over time, the discrepancy between GDP and the more precise methods progressively widens as the number of training epochs increases.
The second plot from the left shows the total privacy leakage computed using the GDP accountant across varying $\delta$ values, from $1\times 10^{-2}$ to $1\times 10^{-6}$. As $\delta$ increases, so does the total privacy leakage ($\varepsilon$), consistent with theoretical expectations. This is because $\delta$ represents the probability of rare but catastrophic privacy failures (e.g., data leakage from physical breaches). Allowing for a higher $\delta$ weakens the privacy guarantee by tolerating a greater failure risk.
@@
Figure 6: Membership risks of NSDEs trained with different configurations. (Top row, left to right) We illustrate the impact of the number of blocks ($n$), the level of stochasticity ($k$), and the step-size ($s$). (Bottom row, left to right) We show the impact of the integration time $T$ and the noise intensity $\sigma$. We use LiRA to assess the membership risks. The results are in CIFAR-10.
@@
Privacy-utility trade-off. The right two plots in Figure 4 illustrate the test accuracy of NSDEs trained with different privacy budgets ($\varepsilon$). We record the test accuracy and privacy leakage at each training epoch for the 16 models used in our evaluation, and we plot them. The second figure from the right shows the trade-offs across different failure probabilities ($\delta$). As $\delta$ increases, the test accuracy of NSDEs increases under the same privacy budget. This is because a larger $\delta$ permits a higher likelihood of privacy leakage. Surprisingly, NSDEs achieve a test accuracy of 80% with $\varepsilon$ = 2 and $\delta$ = $10^{-5}$ under GDP, which demonstrates their ability to maintain high utility even at relatively low privacy budgets. The rightmost plot demonstrates that all three accountants produce similar estimates of privacy leakage and test accuracy. However it is important to note the variability in test accuracy: even under the same privacy budget, e.g., $\varepsilon\!=\!2$, test accuracy ranges from 65–82%. We attribute this variance to the factors like randomness introduced by the diffusion term or the training data stochasticity. In practice, to train NSDEs with high utility we may need to train the same model multiple times, as shown in prior work [47].
5.3 Impact of NSDE Configurations
@@
The top-left plot in Figure 6 illustrates our results. Membership risks associated with NSDE models are consistent when $n\!=\!1$ or $n\!=\!2$. But employing three SDE blocks reduces the risks by a factor of 2, as measured by TPRs at 0.1–1% FPR. We attribute this reduction to poor generalization when $n\!=\!3$. For instance, in CIFAR-10, while NSDEs with $n\in\{1,2\}$ achieve test accuracy of 81.9–82.4%, models with $n\!=\!3$ show significantly lower test accuracy of 74.4%. Interestingly, unlike NODEs, augmenting the number of SDE blocks in NSDEs leads to a corresponding reduction in membership risks.
@@
The top-middle plot illustrates our findings. We show that NSDEs with varying $k$ values effectively mitigate membership risks, as measured by LiRA. With $k$ values up to 0.5, NSDEs exhibit the same level of effectiveness. But when $k$ reaches to 0.6, it completely mitigates the membership risks, bringing the ROC curve below that of a random classifier (a diagonal line). We also examine how each choice of $k$ affects model performance. Increasing the level of stochasticity results in a decrease in model utility. On CIFAR-10, the test accuracy drops from 82.6% to 57.1% as $k$ increases from 0.3 to 0.6. This result implies that the state-of-the-art attack (LiRA) is yet weak so that one can completely mitigate with $k$ that does not significantly degrade model utility.
@@
The top-right plot in the same figure illustrates our findings. Overall, we find no significant interaction between the integration time interval ($T$) and membership risks. All NSDE models consistently show low membership risks. But theoretically, decreasing $T$ will increase the noise intensity ($\sigma$), leading to a smaller privacy leakage ($\varepsilon$) with reduced model utility and membership risks.
@@
The bottom-left plot in Figure 6 summarizes our findings. Overall, we observe only marginal variations in membership risks caused by changes in step size ($1/s$). On CIFAR-10, the membership risks observed at $s\!=\!8$ or $16$ are comparable. However, we find a reduction of the risks with $s\!=\!32$, likely because the model lacks learning detailed dynamics from the training data. Our analysis indicates that this reduction comes at the cost of accuracy. A step size of $32$ limits NSDEs effectiveness in modeling underlying dynamics, and in consequence, they experience substantial performance degradation. NSDEs at this step size achieve an average accuracy of 35.9%, which is 46.5–46.1% lower than that achieved with $s$ of 8 or 16.
@@
The bottom-right plot illustrates our results. Overall, we observe that all the noise intensity values we examine effectively mitigate membership risks, consistent with prior work [6], which shows that even the strongest membership inference attacks can be easily countered by a weak privacy guarantee provided by DP-SGD. We also demonstrate that as the noise intensity gets stronger, NSDE’s effectiveness increases. However, it is important how we configure the noise intensity. For instance, setting a lower noise intensity, such as 1.6, by employing a larger step-size (e.g., 32), may appear to reduce membership riusks in the figure. But this reduction is primarily due to a substantial drop in model utility, as shown in our analysis of different step-sizes.
5.4 NSDE as a Drop-in Replacement Module